How do i solve the equation to see if it is a solution for an ordered pair, I don't understand the substitution or how to substitute what? My equation I have to see if it is a true solution or not to:
(3,2); x+2y=7 What do I use to substitute numbers, my text shows it but it is way too confusing for me to understand.
Thanks,
Millie

Answer by richwmiller(17219)

Your equation is x+2y=7
your point is (3,2) where x=3 and y=2
plug 3 in the equation for x and see if 2 comes up for y
3+2y=7
or
plug 2 in for y and see if x comes up 3
x+2*2=7
or
plug 3 in for x and 2 in for y and see if the result is really equal
3+2*2=7
does the left side really equal the right side?
So there are three similar ways to test it for (3,2)